Problem 6. (10 pts.) Without first computing A-1, find A-1B where 1 0 A= ſi 0 1 2 B = 0 -1 -4 1 -1 1 1 0 1 0
Answer #1

If you have any doubts in the solution please ask me in comments ..

* TO compute AB, without compute A-t. 1 1 - 1 A: B = 0 1 1 1 0 - 0 닛 -1 -4 1 2 211 2012 X3 214 Now assume X- Sot Solution x iX23-133 1 >> 223 = 2 xqui - 2034 1 Rg4 = 2 211 + 3) = 211 = سال) . 212 + 232 = 1213=0 O -> x12 213 + x3 = 1 = Xiutagu - 1 > 2